### v1

# Graph Neural Network-Based Reinforcement Learning for Controlling Biological Networks: The GATTACA Framework

Abstract
Cellular reprogramming, the artificial transformation of one cell type into another, has been attracting increasing research attention due to its therapeutic potential for complex diseases. However, discovering reprogramming strategies through classical wet-lab experiments is hindered by lengthy time commitments and high costs.
In this study, we explore the use of deep reinforcement learning (DRL) to control Boolean network models of complex biological systems, such as gene regulatory networks and signalling pathway networks. We formulate a novel control problem for Boolean network models under the asynchronous update mode in the context of cellular reprogramming. To facilitate scalability, we consider our previously introduced concept of a pseudo-attractor and we improve our procedure for effective identification of pseudo-attractor states. Finally, we devise a computational framework to solve the control problem. To leverage the structure of biological systems, we incorporate graph neural networks with graph convolutions into the artificial neural network approximator for the action-value function learned by the DRL agent. Experiments on a number of large real-world biological networks from literature demonstrate the scalability and effectiveness of our approach.
1 Introduction
Complex diseases remain a major challenge for biomedical research due to the intricate interactions between genes forming gene regulatory networks (GRNs). This network exhibits complex dynamics, emerging from the structure of the network: the state of the genome evolves in time, finally reaching some stable gene expression profiles characterising cell types or, more generally, stable cellular functional states [11].
In some cases, when a GRN is perturbed, it can reach an “unhealthy”, undesired state, that is otherwise inaccessible [5]. An artificial change of the cell fate – called cellular reprogramming aims to solve this issue by steering the GRN back to a “healthy” state. However, identifying effective reprogramming interventions through wet-lab experiments alone is prohibitively expensive and time-consuming. This motivates the development of in-silico methods. To this end, GRNs can be holistically modelled with the formalism of Boolean networks (BNs) originally introduced by S. Kauffman [14]. Its simplicity enables tractable simulations and analysis of non-linear biological dynamics while retaining essential, emergent system behaviour.
To support cellular reprogramming, control strategies must be computed for large-scale BNs, taking into account both network structure and dynamics [9]. Existing state-of-the-art methods, such as CABEAN [30], provide exact algorithms for attractor-based control, but these approaches are limited in scalability due to the exponential growth of the state space and assumptions about network decomposability.
Deep reinforcement learning (DRL) is a promissing alternative for traditional methods for controlling GRNs. It was pioneered in [27] which focused on avoiding undesirable states in terms of steady state probabilities of PBNs. The authors treated gene expression samples as a time-series of experience tuples and used Q-Learning to approximate the optimal policy. To relax the temporal condition, the BOAFQI-Sarsa method was devised in [22]. A further development was mSFQI method proposed in [21], which once again was for control based on the probabilities of gene activity profiles.
The first method that attempted exact control of BNs was the eXtended Classifier System (XCS), which was rule-based reinforcement learning applied to the control of random BNs in [12], and subsequently to the 11-node yeast cell cycle network from [15] in [13].
Scalability is the main concern when devising new methods for control of BNs. All of the methods descirbed above could only control BNs of limited size, with the exact method stopping at 11 nodes. Since deep reinforcement learning (DRL) techniques have proven highly successful in decision problems characterised by large state-action spaces the idea of combining Q-learning with deep learning was conceived. The first ones to realise it were Papagiannis et al. in [23]. They proposed a control method based on deep Q-Learning, namely Double Deep Q-Network (DDQN) with Proritised Experience Replay (PER), initialy introduced in [8]. This method was successful in [24] to control a 20-node synthetic PBN and a 7-node real-world PBN model of melanoma. The same method was used by Acernese et al. in [1] for the control of synchronous PBNs with the goal of driving the networks from a particular state towards a more desirable one, and also in [2] to address the feedback stabilisation problem of 9-gene probabilistic Boolean control networks (PBCNs) modeling apoptosis with the aim of stabilising the system at a given equilibrium point. To tackle the control problem on large networks Moschoyiannis et al. in [19] applied DDQN to control a large synchronous melanoma PBN model of 200 nodes. To make their method scale they limited controll to flipping a single, specific gene, namely pirin, which had already been identified as relevant for the control of PBN models of melanoma in previous studies. Also, the target set of states consisted of exactly half of the entire state space, i.e. the objective was to reach a subset of network states where the WNT5A gene is OFF.
In our previous work [17], we introduced pbn-STAC – a BDQ-based tool for control of asynchronous BNs and probabilistic Boolean networks (PBNs). Therein, we introduced a notions of a pseudo-attractor and a pseudo-attractor state, we devised a pseudo-attractor states identification procedure (PASIP), and we formulated a novel control problem, i.e. the Source-Target Attractor Control, that corresponds with the problem of identifying cellular reprogramming strategies. Our pbn-STAC framework was shown to be able to find optimal and sub-optimal control strategies between source and target attractors for BN and PBN models of sizes up to 33 nodes.
In this study, we improve on our previous work to be able to compute control strategies for very large BN models of biological networks. To this end, we introduce the Graph-based Attractor-Target Control Algorithm (GATTACA), a DRL framework designed to solve a novel target control problem for BN models of GRNs under the asynchronous update mode. As in [17], and unlike other previous approaches, we restrict control actions to be applied only in observable phenotypic cellular states, which are hypothesised to be represented by pseudo-attractors of a BN model. To identify these efficiently, we enhance our earlier pseudo-attractor states identification procedure to handle large and complex networks.
Our method uses a Graph Neural Network (GNN)-based Q-learning agent (ATGC) that learns minimal control strategies to drive the system from a pseudo-attractor state to one aligned with a given target configuration, i.e. specified value for a subset of nodes. The structure of the GRN is directly encoded into the GNN, allowing the model to exploit topological relationships between genes during training.
To assess the scalability potential, we evaluated GATTACA on several large-scale, real-world biological networks, demonstrating its scalability and effectiveness in cases where exact symbolic approaches, such as those implemented in CABEAN, fail. Our contributions represent a step forward in developing predictive, scalable and effective computational methods for cellular reprogramming scenarios.
The paper is organised as follows. The preliminaries are presented in Section 2. Our novel target control problem is the context of cellular reprogramming in formulated in Section 3. Next, a DRL-based framework for controlling large-scale biological networks is developed in Section 4. The framework is evaluated on large models of real-world gene regulatory and signalling networks in accordance with a methodology presented in Section 5. The obtained evaluation results are discussed in Section 6. Finally, the the study is concluded in Section 7.
2 Preliminaries
2.1 Boolean networks
We formally define a Boolean network (BN) as a discrete dynamical system represented by a directed graph $G=(V,E)$, where $V$ denotes the set of nodes representing elements within the system, and $E$ represents the set of directed edges denoting interactions between these elements. Each node $v_{i}$ is associated with a Boolean function $f_{i}$, expressing the state transition rule of the node. These Boolean functions determine the state of each node at time $t+1$, based on the states of its input nodes at time $t$. Mathematically, $f_{i}:\{0,1\}^{k}\rightarrow\{0,1\}$, where $k$ denotes the number of input nodes for node $v_{i}$. The system dynamics evolve iteratively, with the states of the nodes being updated synchronously or asynchronously according to their respective Boolean functions.
A state $x$ of a Boolean network $G=(V,E)$ is a binary vector $x=(x_{1},x_{2},...,x_{n})$, where $n$ is the number of nodes in the network $|V|$, and each $x_{i}$ represents the state of node $v_{i}$ at a particular time step.
A state transition graph (STG) for a Boolean network $G=(V,E)$ is a directed graph that represents the possible state transitions of the network. The nodes of STG correspond to states of $G$, and there exists a directed edge from state $s$ to state $s^{\prime}$ in STG if and only if there exists a Boolean function in $G$ such that applying it to state $s$ results in state $s^{\prime}$. Formally, $STG=(S,\rightarrow)$, where $S$ is the set of all possible states of $G$, and $\rightarrow$ is the set of directed edges representing state transitions.
An attractor of a BN is a bottom strongly connected component (BSCC) in the STG of the network. A fixed-point attractor and a multi-state attractor are BSCCs consisting of a single state or more than one state, respectively.
In the case of the asynchronous update mode, the dynamics of a BN is non-deterministic. It can be modelled as a Markov chain. In this view, an attractor of the BN corresponds to an irreducible subset of states of the Markov chain.
2.2 Pseudo-attractors
It is proven that obtaining the family of all attractors of a BN is NP-hard by itself [3]. This is especially prohibitive in the case of cellular reprogramming considerations, where ultimately large models need to be analysed. However, as we also argued in [17], from the practical point of view, in the context of cellular reprogramming only the frequently revisited states of an attractor are the relevant ones since they correspond to phenotypical cellular states that are observable in the lab. This makes these states ‘recognisable’ for the application of reprogramming interventions in practice in accordance with the control strategy obtained with our framework. With this in mind, we have introduced in [17] the notions of a pseudo-attractor and a pseudo-attractor state in the case of the more general framework of probabilistic Boolean networks (PBNs). Moreover, we proposed the Pseudo-Attractor States Identification Procedure (PASIP) that facilitates the determination pseudo-attractor states in large models. Since in this study our focus is on BN models under the asynchronous update mode, herein we provide a specific formulation of the definition of a pseudo-attractor and a pseudo-attractor state restricted to asynchronous BNs as follows.
Definition 2.1 (Pseudo-attractor).
Let $A$ be an attractor of an asynchronous BN, i.e. an irreducible set of states of the Markov chain underlying the BN. Let $n:=|A|$ be the size of the attractor $A$ and let $\mathbb{P}_{A}$ be the unique stationary probability distribution on $A$. The pseudo-attractor associated with $A$ is the maximal subset $PA\subseteq A$ such that for all $s\in PA$ it holds that $\mathbb{P}_{A}(s)\geq\frac{1}{n}$. The states of a pseudo attractor are referred to as pseudo-attractor states.
Notice that in accordance with the definition, a  pseudo-attractor is a subset of an attractor and pseudo-attractor states are attractor states.
As argued in [17], the correctness of the definition is guaranteed by the fact that the state space of the underlying Markov chain of a BN is finite and that the Markov chain restricted to the attractor states is irreducible. It is a well known fact that all states of a finite and irreducible Markov chain are positive recurrent. In consequence, the attractor-restricted Markov chain has a unique stationary distribution.
2.3 Reinforcement learning
The Attractor-Target Control can be interpreted as a Markov Decision Process (MDP) and it can be solved using reinforcement learning (RL). RL solves sequential decision problems by optimizing a cumulative reward function.
The RL problem is meant to be a direct framing of the task of learning from interaction to achieve a goal. The decision-maker and learner is referred to as the agent. At each time step $t$, the agent interacts with the environment which comprises everything outside the agent by implementing a mapping $\pi:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ from states to probabilities of selecting an action $A_{t}$ in state $S_{t}$. This mapping is referred to as the agent’s policy, where $\pi(a\mid s)$ is the probability that $A_{t}=a$ given that $S_{t}=s$.
Within the MDP-based framework, the RL problem is formalised as the optimal control of MDPs. RL methods specify how the agent changes its policy as a result of its experiences with the goal of maximising the expected discounted return:
$G_{t}\triangleq\sum_{k=0}^{\infty}\gamma^{k}R_{t+k+1},$
where $0\leq\gamma\leq 1$ is the discount factor.
Nearly all RL algorithms involve estimating value functions. For MDPs, two such functions are commonly considered. First, the state-value function $v_{\pi}(s)$ defined as
$v_{\pi}(s)\triangleq\mathbb{E}_{\pi}[G_{t}\mid S_{t}=s]=\mathbb{E}_{\pi}\left[\sum_{k=0}^{\infty}\gamma^{k}R_{t+k+1}\bigm{|}S_{t}=s\right]\!,$
where $\mathbb{E}_{\pi}[\cdot]$ is the expected value of a random variable given that the agent follows policy $\pi$. Second, the action-value function $q_{\pi}(s,a)$ is defined as the expected return starting from $s$, taking the action $a$, and thereafter following policy $\pi$:
$q_{\pi}(s,a)$ $\triangleq\mathbb{E}_{\pi}[G_{t}\mid S_{t}=s,A_{t}=a]$
$=\mathbb{E}_{\pi}\left[\sum_{k=0}^{\infty}\gamma^{k}R_{t+k+1}\biggm{|}S_{t}=s,A_{t}=a\right]\!.$
The value functions satisfy particular recursive relationships expressed in the form of Bellman equations, i.e. one for $v_{\pi}$:
$v_{\pi}(s)=\sum_{a\in\mathcal{A}}\pi(s,a)\sum_{s^{\prime}\in\mathcal{S}}P_{a}(s,s^{\prime})\Bigl{[}R_{a}(s,s^{\prime})+\gamma v_{\pi}(s^{\prime})\Bigr{]}$ (1)
and one for $q_{\pi}$:
$q_{\pi}(s,a)=\sum_{s^{\prime}\in\mathcal{S}}P_{a}(s,s^{\prime})\Bigl{[}R_{a}(s,s^{\prime})+\gamma\sum_{a^{\prime}\in\mathcal{A}}\pi(s^{\prime},a^{\prime})q_{\pi}(s^{\prime},a^{\prime})\Bigr{]}.$ (2)
Value functions define a partial order over policies, that is, $\pi\geq\pi^{\prime}$ if and only if $v_{\pi}(s)\geq v_{\pi}^{\prime}(s)$ for all $s\in\mathcal{S}$. There always exists at least one policy which is better than or equal to all other policies [31]. Such policy is referred to as an optimal policy and all optimal policies are denoted by $\pi_{*}$. Optimal policies share the same optimal state-value function, denoted $v_{*}$, and defined as
$v_{*}(s)\triangleq\max_{\pi}v_{\pi}(s),$
for all $s\in\mathcal{S}$. They also share the same optimal action-value function, denoted $q_{*}$, which is defined as
$q_{*}(s,a)\triangleq\max_{\pi}q_{\pi}(s,a),$
for all $s\in\mathcal{S}$. It holds that
$q_{*}(s,a)=\mathbb{E}_{\pi_{*}}\bigl{[}R_{t+1}+\gamma v_{*}(S_{t+1})\bigm{|}S_{t}=a,A_{t}=a\bigr{]}.$
The optimal value functions satisfy the Bellman optimality equations, i.e.
$v_{*}(s)$ $=\max_{a\in\mathcal{A}}\mathbb{E}\bigl{[}R_{t+1}+\gamma v_{*}(S_{t+1})\bigm{|}S_{t}=s,A_{t}=a\bigr{]}$
$=\max_{a\in\mathcal{A}}\sum_{s^{\prime}\in\mathcal{S}}P_{a}(s,s^{\prime})\bigl{[}R_{a}(s,s^{\prime})+\gamma v_{*}(s^{\prime})\bigr{]}$
and
$q_{*}(s,a)$ $=\mathbb{E}\bigl{[}R_{t+1}+\gamma\max_{a^{\prime}\in\mathcal{A}}q_{*}(S_{t+1},a^{\prime})\bigm{|}S_{t}=s,A_{t}=a\bigr{]}$
$=\sum_{s^{\prime}\in\mathcal{S}}P_{a}(s,s^{\prime})\bigl{[}R_{a}(s,s^{\prime})+\gamma\max_{a^{\prime}\in\mathcal{A}}q_{*}(s^{\prime},a^{\prime})\bigr{]},$
where the expectations are taken without reference to any specific policy since the value functions are optimal.
For finite MDPs, the Bellman optimality equations have unique solutions independent of the policy. If the dynamics of the MDP in known, i.e. the $P$ and $R$ functions, one can solve the Bellman optimality equations. Once the optimal state-value function $v_{*}$ is known, any policy that is greedy with respect to $v_{*}$ is an optimal policy.
Optimal policies can be computed with dynamic programming (DP) algorithms, such as policy iteration or value iteration, given a perfect model of the environment in the form of an MDP. DP algorithms are obtained by turning Bellman equations into update rules that improve the approximations of the desired value functions. However, the utility of DP algorithms is limited due to the perfect model assumption and its computational expense. Therefore, for large state space environments or cases there the model of the environment is unavailable, Monte Carlo methods or temporal-difference learning (TD) methods are used, see [31].
2.4 Q-learning
Q-learning is a TD control algorithm and is one of the most popular model-free methods for solving RL problems. It is used to learn the optimal action-selection policy. The one-step Q-learning, which is the simplest form of Q-learning, is defined by the update rule
$Q(S_{t},A_{t})\leftarrow Q(S_{t},A_{t})+\alpha\bigl{[}R_{t+1}+\gamma\max_{a\in\mathcal{A}}Q(S_{t+1},a)-Q(S_{t},A_{t})\bigr{]}\!,$
where $0<\alpha\leq 1$ is the learning rate, which determines how fast past experiences are forgotten by the agent, and the learned action-value function $Q$ directly approximates $q_{*}$. The algorithm can be outlined as follows. First, initialise $Q(s,a)$ for all $(s,a)\in\mathcal{S}\times\mathcal{A}$ and start the environment at state $S_{0}$. For time steps $t=0,1,2,\ldots$, choose action $A_{t}$ according to the $\epsilon$-greedy policy, i.e.
$A_{t}\leftarrow\begin{cases}\textrm{argmax}_{a\in\mathcal{A}}Q(S_{t},a)&\textrm{with probability }1-\epsilon\\ \textrm{uniformy random action in }\mathcal{A}&\textrm{with probability }\epsilon,\\ \end{cases}$
and take action $A_{t}$ in the environment. In result, the state is changed from $S_{t}$ to $S_{t+1}\sim P_{a}(s,\cdot)$. Observe $S_{t+1}$ and $R_{t+1}$. Finally, revise the action-value function $Q$ at state-action $(S_{t},A_{t})$ in accordance with the above update rule.
The Q-learning algorithm iteratively updates the Q-values based on observed transitions, gradually refining its estimates of the optimal action-value function. All that is required to assure convergence to $q_{*}$ is that all $(s,a)\in\mathcal{S}\times\mathcal{A}$ pairs continue to be updated. Notice that the convergence is guaranteed independently of the policy being followed. This property renders Q-learning an off-policy TD control algorithm and enables the use of the $\epsilon$-greedy policy, which is a simple technique for handling the exploration-exploitation tradeoff, to ensure the convergence. Once learned, the Q-values are used to provide the agent with the optimal policy for selecting actions in each state.
2.5 Deep reinforcement learning
The Q-learning algorithm has proved successful for controlling small-scale environments. However, in the case of large state-action spaces, Q-learning cannot learn all action values in all states separately and can fail to converge to an optimal solution. To address this issue, the idea was conceived to learn a parameterised value function $Q(a,s;\theta)$ by updating the parameters $\theta$ in accordance with the standard Q-learning-derived rule:
$\theta_{t+1}\leftarrow\theta_{t}+\alpha\Big{[}R_{t+1}+\gamma\max_{a}Q(S_{t+1},a;\theta_{t})$
$-Q(S_{t},A_{t};\theta_{t})\Big{]}\nabla_{\theta_{t}}Q(S_{t},A_{t};\theta_{t}),$
resembling stochastic gradient descent. It was shown that as the agent explores the environment, the parameterized value function converges to the true $Q$-function, see, e.g., [31].
The above idea is realized in the deep RL (DRL) framework, which combines classical RL with artificial neural network (ANN) function approximation. In DRL, a function approximator is trained to estimate the Q-values. The approximator is a deep Q-network agent (DQN), i.e. a multi-layer ANN which for a given observed state of the environment $s$ outputs a vector of action values $Q(s,\cdot;\theta)$, where $\theta$ are the parameters of DQN. The aim of DQN is to approximate $q_{*}$ by iteratively minimising a sequence of loss functions $L_{t}(\theta_{t})$:
$L_{t}(\theta_{t})=\left(Y_{t}^{\textrm{DQN}}-Q(S_{t},A_{t};\theta_{t})\right)^{2},$
where $Y_{t}^{\textrm{DQN}}$ is the DQN target defined as
$Y_{t}^{\textrm{DQN}}\triangleq R_{t+1}+\gamma\max_{a}Q(S_{t+1},a;\theta_{t}^{-}),$
while $\theta_{t}^{-}$ and $\theta_{t}$ are the parameters of the target and online networks, respectively. The target network is the same as the on-line network, but its parameter values are copied from the online network every $k$ steps and kept fixed at other steps.
2.5.1 Double Deep Q-Network
In the study of [18], the Double Deep Q-Network (DDQN) [34] was used to control BNs. Both standard Q-learning and DQN uses the same values both to select and to evaluate an action, which can lead to selection based on overestimated values [34]. DDQN combines Double Q-learning [34] with Deep Q-learning to address this issue and uses the following DDQN target In the study  [18], the Double Deep Q-Network (DDQN) from  [34] was used to control Boolean Networks. Both standard Q-learning and Deep Q-Network (DQN) use the same values for both selecting and evaluating actions, which can lead to selecting actions based on overestimated values, as noted by the authors. To address this issue, DDQN introduces Double Q-learning with separete networks for predicting actions and their values. It uses the following DDQN target:
$Y_{t}^{\textrm{DDQN}}\triangleq R_{t+1}+\gamma Q(S_{t+1},\textrm{argmax}_{a}Q(S_{t+1},a;\theta_{t});\theta_{t}^{\prime}).$
The second set of parameters, $\theta_{t}^{\prime}$, enables fairly evaluation of the greedy policy. The DDQN parameters are updated by symmetrically switching the roles of $\theta_{t}$ and $\theta_{t}^{\prime}$.
2.5.2 Branching Dueling Q-Network
In our study, to allow for simultaneous perturbation of a combination of genes within a DRL action, we replace the original DDQN architecture of [18] with the Branching Dueling Q-Network (BDQ) architecture [33]. BDQs are designed to address complex and high-dimensional action spaces; they enhance the scalability and sample efficiency of RL algorithms in complex scenarios. BDQs can be thought of as an adaptation of the dueling network architecture [36] into the action branching architecture. A BDQ uses two separate ANNs, i.e. the target network for evaluation and the controller network for selection of actions. Furthermore, instead of using a single output layer for all actions, BDQ introduces multiple branches, each responsible for a specific subset of actions. This decomposition allows for parallel processing and shared representations, making it more efficient for the agent to learn and execute tasks with a diverse action space. BDQs avoid overestimating Q-values, can more rapidly identify action redundancies, and generalise more efficiently by learning a general Q-value that is shared across many similar actions.
2.6 Graph Neural Networks
Graph Neural Networks (GNNs) have gained significant attention in recent years due to their ability to model and analyse graph-structured data. Graphs are powerful generic mathematical models that capture relationships and dependencies between entities in various domains, including social networks, biological systems, and recommendation engines. The main idea underlying GNNs is the use of pairwise message passing, i.e. graph nodes iteratively update their representations by exchanging information with their direct neighbours. In this study we focus on a specific type of GNNs, i.e. Graph Convolutional Networks.
2.6.1 Graph Convolutional Operation
Graph Convolutional Networks (GCNs) are a subclass of GNNs and can be considered a generalization of classical convolutional neural networks (CNN) to graph-structured data. The essence of GCNs lies in their ability to perform graph convolutions, a key operation that allows them to learn and propagate information across the nodes of a graph. Graph convolution is inspired by traditional convolutional operations in image processing. While convolutional layers in traditional neural networks operate on grid-structured data, graph convolutional layers extend this concept to non-grid structures. The goal is to capture and aggregate information from neighbouring nodes, enabling the model to understand the relational structure within the graph. The information propagation happens through kernels – that is functions similar to kernels in CNNs but adapted to the graph’s structure.
In our implementation, a graph convolutional layer takes the values of each input node from the previous layer and produces the corresponding output vectors for each node. The convolutional layer implements the function:
$x^{\prime}_{v}=\sum_{u\in N(v)}h_{\Theta}(v||u-v),$
where $v$ is the current node, $N(v)$ is the set of neighbours of $v$, and $h$ is the graph convolutional kernel with learnable parameters $\Theta$. In our case $h$ consists of a neural network with a 64-neuron deep layer and ReLu activations. Here $h(x||y)$ means conditioning-by-intervention, that is the distribution of $x$ with $y$ set to some fixed value. It’s role is to reduce spurious correlations present in the graph; for more details, we refer to [35]. Our network consists of three graph convolutional layers. The output of graph convolutions is then passed through a three-hidden-layer MLP with 1024, 512, and 256 neurons per layer. Each of the attention branches, as well as the state value layer, is an MLP with (256, 512, 512) neurons. We expect that our network will learn interactions between genes, and create an embedding of a discreet graph into continuos space, which is more suitable for machine learning algorithms.
3 Control problem formulation
We proceed to formulate a novel version of the target control problem tailored for cellular reprogramming, understood as an artificial change of the identity or functional state of a cell by manipulating its gene expression profile.
Often biologists are interested not in a single state but a set of states in which a specified subset of genes attains certain configuration of expression levels which afterwards remains constant. For example, they may be interested in achieving any state in which apoptosis occurs. Therefore, we assume that the target is defined as a combination of desired values for a subset of genes. With this, we can proceed with the definition of an attractor-target control strategy.
source attractor state target1target2target3controlcontrolcontroltarget basinintermediate basinintermediate basin target configuration
Figure 1: Schematic illustration of the concept behind Attractor-Target Control. The target state can be any of the attractor states that are aligned with the target configuration. Transitions that follow the original network dynamics are shown as dashed arrows. Red lightnings indicate interventions. The depicted control strategy is of length 3.
Definition 3.1 (Attractor-Target Control Strategy).
Given a BN, a source (pseudo-)attractor state, and a specified target configuration for a subset of genes, an attractor-target control strategy is a sequence of interventions that guides the BN dynamics from the source (pseudo-)attractor state to some (pseudo-)attractor state that aligns with the desired gene configuration. Interventions are understood as simultaneous flips (perturbations) of values for a subset of genes in a particular network state and their application is limited to (pseudo-)attractor states.
Furthermore, length of a control strategy is defined as the number of interventions in the control sequence. We refer to a target control strategy of the shortest length as the minimal attractor-target control strategy.
The concept of the attractor-target control strategy is schematically illustrated in Figure 1. Notice that, according to the above definition, we aim to guide the network dynamics toward any state within a subset of (pseudo-)attractor states that match the target configuration, rather than a specific one. This definition is more general than the definition of the ‘Attractor-Based Control Strategy’ introduced in [17], where the system is expected to reach a specific attractor. The current definition of the ‘Attractor-Target Control Strategy’ aligns better with contemporary biological practice.
We now define the attractor-target control problem, which is a variation of the ‘target-based sequential instantaneous control (ASI) problem defined for BNs in [28]. An exact ‘divide-and-conquer’-type algorithm for solving the ASI problem for BNs is implemented in the software tool CABEAN [30], which we use as a benchmark in the evaluation of our proposed approach.
Definition 3.2 (Attractor-Target Control).
Given a BN, a source (pseudo-)attractor, and a target configuration for a subset of genes, find a minimal attractor-target control strategy.
4 Graph-based Attractor-Target Control Algorithm
We propose herein a Graph-based Attractor-Target Control Algorithm (GATTACA) – a GNN-based DRL framework for the attractor-target control problem. We focus on cellular phenotypic functional states, which are observable in experimental practice, and thus allow the DRL agent, which we refer to as the Attractor-Target Graph-based Controller (ATGC), to take actions only in (pseudo-)attractor states in agreement with our control problem formulation in Section 3. Since the identification of all attractor states is an NP-hard problem in itself, we further develop the pseudo-attractor identification procedure (PASIP), which we previously established in [17]. The source of GATTACA and the trained models are available at [37].
4.1 Improved PASIP
We observed that the original PASIP method in [17] can be impeded by the presence of large attractors in a BN model. For example, the T-LGL network of [38] has an attractor with more than 1000 states. Of course, if the existence of a large attractor is foreseen and a rough estimate of its size is available, one could adjust the PASIP hyperparameters appropriately. However, this would imply the necessity of longer simulations of the BN model, which are computationally expensive. If the existence of a large attractor is not anticipated and the hyper-parameters are not adjusted, the original PASIP is prone to getting trapped in such an attractor while waiting for one of the attractor states to demonstrate a pronounced frequency of re-visits. To handle such cases effectively, we modify PASIP by adding additional checkpoints that enable the identification of a representative state within a large attractor. The improved PASIP procedure (iPASIP) consists of steps executed in two phases: (I) environment pre-processing and (II) DRL agent training. It extends the original PASIP from [17] by incorporating Steps I-2 and II-3 as the additional checkpoints.
  1. I-1
    During environment pre-processing, $k$ PBN simulations are started from randomly selected initial states. Each PBN simulation runs for initial $n_{0}=200$ time steps, which are discarded as the so-called burn-in period. Then, the simulation continues for $n_{1}=1000$ time steps during which the visits to individual states are counted. All states in which at least $5\%$ of the simulation time $n_{1}$ is spent are added to the list of pseudo-attractor states.
  2. I-2
    If at any point more than 1,000,000 steps are taken without encountering a dominating state (one revisited at least 5% of the time), it is assumed that the dynamics have become trapped in a large attractor. The most frequently encountered state is then added to the list of pseudo-attractor states as a representative of this plausible attractor.
  3. II-1
    The simulation of the PBN environment may enter a fix-point attractor not detected in Step I. If the simulation gets stuck in a particular state for $n_{2}=1000$ steps, the state is added to the list of pseudo-attractor states.
  4. II-2
    During training, the simulation of the PBN environment may enter a multi-state attractor that was not detected in Step I. To account for this, a history of the most recently visited states is maintained. When the history buffer reaches a size of $n_{3}=10000$ items, revisits to each state are counted and any state revisited more than 15% of the time is added to the list of pseudo-attractor states. If no such state exists, the history buffer is cleared and the procedure continues for another $n_{3}$ time steps. Otherwise, the history information is discarded.
  5. II-3
    As in I-2, if at any point more than 1,000,000 steps are taken without encountering a dominating state (one revisited at least 5% of the time), it is assumed that the dynamics have become trapped in a large attractor and a representative of this plausible attractor is added to the list of pseudo-attractor states.
We notice that while the new checkpoints may clearly introduce false-positive pseudo-attractor states, this occurs rarely. Importantly, the representative states of plausible large attractors stand out due to being frequently revisited. Even if these states are not exact pseudo-attractor states, to a large extent they resemble them. Therefore, given their rarity, adding them to the list of pseudo-attractor states is acceptable. This is confirmed by the results of our case studies, presented and discussed in Section 6.
Our control learning algorithm can handle cases where the number of pseudo-attractors is greater than 25,000. Moreover, we conducted experiments in which the number of pseudo-attractor states was overestimated by a factor of 1000 and ATGC was still able to identify effective control strategies (data not shown). Those two facts prove that the ATGC is resilient to both high attractor counts and spurious pseudo-attractor states.
Since iPASIP method can introduce spurious pseudo-attractor states, we will henceforth use the term ‘pseudo-attractor state’ in a broader sense: it will refer to all states returned by iPASIP rather than only those defined in Definition 2.1.
4.2 Attractor-Target Graph-based Controller (ATGC)
In our training framework, we specify a source pseudo-attractor state and a target configuration that specifies the desired values for a subset of genes. Then, we let the DRL agent perform actions, i.e. interventions consisting of simultaneous perturbation of up to five genes once a pseudo-attractor state different from the target is reached. After an action, we let the environment evolve in accordance with its original dynamics into the next pseudo-attractor state. This models the biological process of transforming a cell from one phenotypic functional state into another by conducting a series of in vivo reprogramming interventions.
Figure 2: Schematic representation of the ATGC DRL agent of GATTACA.
We define a new with respect to the one in [17] DRL reward function as
$R_{a}(s,s^{\prime})=21+100*\mathbbm{1}_{TA}(s^{\prime})-|a|,$
where $s$ and $s^{\prime}$ are the current and next pseudo-attractor states, respectively, $\mathbbm{1}_{TA}$ is an indicator function of the target configuration, and $|a|$ is the number of genes perturbed by applying action $a$.
This reward scheme is balanced to punish the model for perturbing many genes in a single action while providing enough reward to allow to still reach the target configuration by the means of a sequence of actions. Our previous experiments showed that it is important that the reward function maintains a consistent sign during training (in this particular case it is designed to remain non-negative, which is ensured by the term of 21). For the loss function, We use the Double Deep Q Network (DDQN) loss function as in [18] and [17], that is the Mean Squared Error (MSE) between the predicted Q-values and the target Q-values, calculated using equation:
$\mathcal{L}(\theta)=\mathbb{E}\left[\left(q_{\pi}(s,a)-target(s,a)\right)^{2}\right],$
where $target$ is the delayed copy of the $q_{\pi}$ network.
Our DRL agent is trained in episodes. An episode is a sequence of states, actions, and rewards that starts in the specified source state and ends in any of the pseudo-attractor states that comply with the target configuration. It represents a complete trajectory of the agent’s interaction with the environment.
For each episode, we randomly choose a source (pseudo-)attractor state. We make sure that this state fulfils the imposed initial conditions and does not satisfy the target configuration. Then we let our agent apply its strategy and terminate each episode after 100 unsuccessful actions. If 100 actions are not enough to reach the target configuration, the respective episode is considered unsuccessful and terminated.
5 Experiments
We evaluate how GATTACA scales with the size of the BN model. We focus on large, verified models taken from the literature for well-known real-world biological networks.
5.1 Case Studies: Models and Their Target Configurations
To evaluate the performance of our approach, we select a variety of models from the literature, ranging in size from 35 nodes (the Bladder model) to 188 nodes (the CD4+ model). To ensure that the target configurations used in our evaluation are both realistic and of significance to biologists, we focus on the configurations considered in the original publications where these models were introduced. The models and their corresponding target configurations are listed in Table 1. Below, we briefly present each of them.
Model BD Id Size Target Configuration
Bladder 183 35 Apoptosis_b1=1 or Apoptosis_b2=1
MAPK 70 53 Apoptosis=1
T-LGL 14 61 Apoptosis=1
Bortezomib 59 67 JNK=p21=Cas3=1
T-diff - 68 / 71 TBET=RORGT=FOXP3=1, GATA3=0
ABA 84 81 Closure=1
MCF-7 232 117 CASP=E2F1=1
CD4+ 50 188 GATA3=1
Table 1: BN models of real-world biological networks with their target configurations used for the performance evaluation of GATTACA. The BD Id column provides model identification numbers in the BioDivine repository [25].
Model Size (Inputs) Environmental Condition
Bladder 35 (4) EGFR_stimulus = 0, DNAdamage = *, FGFR3_stimulus = *, GrowthInhibitors = 1
MAPK 53 (4) EGFR_stimulus = 0, FGFR3_stimulus = 0, Growth_Arrest = *, DNA_damage = *
T-LGL 61 (7) CD45 = *, IFN = *, IL15 = *, PDGF = *, Stimuli = *, Stimuli2 = *, TAX = *
Bortezomib 67 (5) XX = 0, SHP1 = 0, TNFAR = 0, Bort = 0, TNFA = 0
T-diff 68 / 71 (24) APC = 1, IFNB_e = 0, IFNG_e = 0, IL2_e = 1, IL4_e = 0, IL6_e = 0, IL10_e = 0, IL12_e = 0, IL15_e = 0, IL21_e = 0, IL23_e = 0, IL27_e = 0, TGFB_e = 1, IFNGR1 = 1, IFNGR2 = 1, GP130 = 1, IL6RA = 1, CGC = 1, IL12RB2 = 1, IL10RB = 1, IL10RA = 1, IL15RA = 1, IL2RB = 1, IL27RA = 1
ABA 81 (23) PC = *, PtdInsP3 = *, DAGK = *, CPK23 = *, NAD = *, Sph = *, SCAB1 = *, Nitrite = *, ERA1 = *, ABH1 = *, SPP1 = *, GEF1_4_10 = *, ARP_Complex = *, MRP5 = *, RCN1 = *, NtSyp121 = *, ABA = *, NADPH = *, GCR1 = *, PtdInsP4 = *, C1_2 = *, GTP = *, CPK6 = *
MCF-7 117 (8) DLL_i = 1, EGF = 1, ES = 1, IGF1 = 1, INS = 1, NRG1 = 1, PG = 1, WNT1 = 1
CD4+ 188 (34) IL27RA = 0, IL27_e = 0, GP130 = 0, Galpha_QL = 0, IL2RB = 0, CGC = 0, Galpha_iL = 0, MHC_II = 0, APC = 0, IL18_e = 0, IL9_e = 0, IFNB_e = 0, ECM = 0, IL21_e = 0, alpha_13L = 0, IL10RA = 0, IL10RB = 0, IL10_e = 0, IL15_e = 0, B7 = 0, IFNGR1 = 0, IFNGR2 = 0, IFNG_e = 0, CAV1_ACTIVATOR = *, GalphaS_L = 0, IL4_e = 0, IL6_e = 0, IL6RA = 0, TGFB_e = 0, IL22_e = 0, IL2_e = 0, IL23_e = 0, IL15RA = 0, IL12_e = 0
Table 2: Environmental conditions for the BN models of real-world biological networks used in the evaluation of GATTACA. Unspecified node values are indicated with *.
Bladder (Bladder)
The logical model for the prediction of bladder cancer cells to become invasive was originally introduced in [26] with the purpose of understanding how genetic alterations combine to promote cancer tumorigenesis. Specifically, the focus is on the relationships between the alterations, i.e. mutual exclusivity and co-occurrence. Using statistical analysis on tumour datasets and a logical modelling approach, the study explores patterns of genetic alterations that affect growth factor receptor signalling, cell cycle regulation, and apoptosis. The model confirms conditions where genetic alterations drive tumour progression and suggests that additional mutations, such as PIK3CA or p21CIP deletion, may enhance invasiveness.
MAPK cancer cell fate (MAPK)
The Mitogen-Activated Protein Kinase (MAPK) network is a system of tightly interconnected signalling pathways involved in diverse cellular processes, such as cell cycle, differentiation, survival, and apoptosis. In [10], a comprehensive and generic reaction map of the MAPK signalling network is constructed based on an extensive analysis of published data. To investigate MAPK responses to different stimuli and understand their involvement in cell fate decisions in urinary bladder cancer, specifically the balance between cell proliferation and cell death, a logical model comprising 53 nodes is constructed, incorporating the most crucial components and interactions. The Boolean model is verified to be globally consistent with published data and it is capable of recapitulating salient dynamical properties of the complex biological network.
T-LGL survival network (T-LGL)
T cell large granular lymphocyte (T-LGL) leukemia features a clonal expansion of antigen-primed, competent, cytotoxic T lymphocytes (CTL). To understand the signalling components that are crucial for the survival of CTL in T-LGL leukemia, a T-LGL survival signalling network is constructed in [38]. The network integrates the signalling pathways involved in normal CTL activation and the known deregulations of survival signalling in leukemic T-LGL. The network is translated into a BN model that describes the signalling involved in maintaining the long-term survival of competent CTL in humans. The BN model is capable of reproducing a clinically relevant complex process. In particular, it predicts that the persistence of IL-15 and PDGF is sufficient to reproduce all known deregulations in leukemic T-LGL. For further details, we refer to [38].
ABA-induced stomatal closure (ABA)
A vital process for drought tolerance in plants is examined in [4]. To reduce water loss during drought stress conditions, the phytohormone abscisic acid (ABA) triggers the closure of stomata, i.e. microscopic pores on leaf surfaces through which water loss and carbon dioxide uptake occur. Based on an extensive review of the literature on ABA signal transduction responses underlying stomatal closure, a network of 84 nodes and 156 edges is constructed. This network is then translated into a BN model which predictions agree with 82 out of 99 considered experiments. The model reveals nodes that could be controlled to influence stomatal closure and provides a substantial number of novel predictions that could be subject to wet-lab validation.
MCF-7 (MCF-7)
Cell growth and proliferation in breast cancer incorporating common signalling pathways is analysed in [32]. To account for the fact that in different cell lines the same combination of drugs can produce different synergistic effects leading to heterogenous responses, the generic model is tailored to represent five breast cancer cell lines. In particular, by integrating information on cell line specific mutations, gene expression profiles, and observed behaviours from drug-perturbed experiments, a BN model for the MCF-7 cell line is proposed. The model predictions on cell-line specific protein activities and drug-response behaviours are in agreement with experimental data.
T-diff (T-diff)
A comprehensive 71-node Boolean network model was developed to represent the regulatory and signaling pathways controlling CD4+ T helper (Th) cell differentiation [20]. The model captures key transcription factors, cytokine receptors, and intracellular signaling components involved in Th lineage specification. Simulations revealed the emergence of stable states corresponding to canonical Th1, Th2, Th17, and Treg subsets, as well as hybrid cell types that co-express lineage-specific markers, reflecting the context-dependent plasticity of Th cells. Environmental inputs, such as cytokine concentrations, were shown to modulate the attainable cell fates, emphasizing the model’s ability to integrate external cues. A reduced 68-node version of the original network was subsequently used as a case study in [29], focusing on specific environmental conditions to investigate dynamic transitions between Th phenotypes. This work demonstrated the model’s applicability to studying immune cell adaptation in diverse microenvironments and highlighted potential regulatory checkpoints for therapeutic intervention.
Bortezomib
[…]
CD4+
Caveolin-1 (CAV1) is a vital scaffold protein heterogeneously expressed in both healthy and malignant tissue. In [7], the role of overexpressed CAV1 in T-cell leukemia is studied. In particular, a 188-node BN model of a CD4+ immune effector T-cell is constructed. The model incorporates various signalling pathways and corresponding protein-to-protein, protein-phosphorylation, and kinase interactions to mimic cellular dynamics and molecular signalling under healthy and immunocompromised conditions. The model was used to predict and examine in silico the heterogeneous effects and mechanisms of CAV1. Specifically, to verify the hypothesis that the role of CAV1 in immune synapse formation contributes to immune regulation during leukemic progression. It was verified that the local interactions built into the BN model were able to accurately mimic complex phenomena observed in the laboratory.
5.2 Performance Evaluation Methodology
We evaluated the performance of GATTACA in solving the attractor-target control problem formulated in Section 3 on asynchronous BN models of real biological systems in a wide range of sizes using case studies described in Section 5.1.
In logical models, input nodes typically define specific environmental conditions and are kept fixed. Therefore, when going from one attractor in a given environmental condition to another attractor in a different condition, the input nodes always need to be set accordingly. Hence, they are not very interesting from the control problem point of view. That is why for the Bortezomib, T-diff, MCF-7, and CD4+ models we consider a specific, biologically relevant environmental condition taken from the respective publication in which a particular model was originally introduced. However, for the smaller Bladder and MAPK models, we fix only a subset of input nodes, which in both cases leads to four environmental conditions being considered at once. This is due to the fact that in the majority of cases where all input nodes are set to a specific environmental condition, the models manifest either a single attractor or no or all attractors aligned with the target configuration. Hence, the control problem has either no solution or trivial ones, where no control action is needed. Therefore, we allow more than one fully specified environmental condition to be considered at once by allowing any value for some of the input nodes.
For the T-LGL and ABA models, no environmental conditions were specified in the publications in which they were introduced. Therefore, we consider all possible environmental conditions for these two models.
Since we want ATGC of GATTACA to learn the relationship between the target genes and other genes, we let it perturb all nodes (including input nodes) except the ones specified in the target configuration, which we refer to as target genes. Beyond this, we do not impose any other restrictions on genes that can be perturbed: we allow full flexibility in selecting genes for interventions to ensure a fair comparison with the CABEAN benchmark and to demonstrate the potential of GATTACA to identify optimal or suboptimal strategies in large networks. However, we emphasize that our framework can be easily modified to incorporate previous biological knowledge by restricting perturbations to a predefined subset of genes. In our experiments, we specify the target genes and their configurations based on the literature from which the models were sourced. The details are presented in Section 5.1.
Furthermore, the more genes that need to be perturbed in a single intervention, the more challenging and expensive the experimental procedure becomes. As a result, the in vitro implementation of the predicted control strategy may be unfeasible. Hence, to make our approach as cost-effective as possible, we decide to limit the number of nodes that can be perturbed simultaneously in a single intervention. This number is a parameter of our method and we set it to five in our experiments. This setting suffices to compute successful control strategies for all of our case studies. Nevertheless, the parameter value can be adjusted to particular needs as required.
ATGC is trained to minimise the number of control actions required to drive a BN model from a randomly chosen source attractor to any attractor in which the given target configuration is realised. Specifically, for a specified environmental condition, a training episode begins by randomly selecting one of the pseudo-attractor states identified so far by iPASIP that does not align with the given target configuration. Throughout each episode, we track the number of control actions taken, referred to as the episode control length, as ATGC guides the network dynamics toward any pseudo-attractor state aligned with the given target configuration. ATGC can take up to 100 control actions per episode. If it fails to control the BN model within this limit, the episode is considered unsuccessful, indicating that no valid control strategy was found.
To evaluate the performance of GATTACA on a particular BN, we recover control strategies by simulating the environment from a randomly selected source pseudo-attractor state that is not aligned with the target configuration. At each intermediate pseudo-attractor state, a control action predicted by the trained DRL agent is applied. For each recovered control strategy, we record its length and the information whether the target configuration is reached. Due to the nondeterministic dynamics of the environment introduced by the asynchronous update mode, results may vary between runs. To address this issue, we recover a control strategy 10 times for each of the randomly selected source pseudo-attractor states. For a given BN model, we report the average length of successful control strategies.
We compare the results of GATTACA with those obtained using the state-of-the-art exact attractor-based sequential instantaneous source-target control (ASI) algorithm introduced in [16] and implemented in the CABEAN software tool [30]. The exact algorithm follows the divide-and-conquer approach based on SCC-decomposition of BN structure graph and it leverages symbolic techniques based on Binary Decision Diagrams (BDDs). Amongst the control algorithms provided by CABEAN, we selected ASI as the best match for the control problem in this study.
For a fair comparison, we set in CABEAN the initial values of the input nodes of a BN model in accordance with a particular environmental condition. CABEAN requires at least some of the input nodes to have fixed values. Therefore, in the case of partially specified environmental conditions, i.e. for models Bladder, MAPK, and CD4+, it is possible to consider all of the corresponding fully specified environmental conditions at once as is the case for GATTACA.
However, in the case of all environmental conditions considered, e.g. the T-LGL model, fully specified environmental conditions are enumerated and considered one by one in CABEAN. For each such condition, all attractors are computed with CABEAN and divided into source attractors, i.e. those not aligned with the target configuration, and target attractors, i.e. those aligned with the target configuration. Then, for each source attractor, the shortest path to any of the target attractors is computed with CABEAN. If there is no available target attractor for a given environmental condition, we check all other environmental conditions for a target attractor. We first switch the environmental condition, and then (if necessary) use CABEAN to control the network within the new environmental condition. The environmental condition switch adds 1 to the control path length in CABEAN. We take the minimum length over all other environmental conditions containing a target attractor. We provide an example of this scenario in Appendix B. This behaviour is mimicked by GATTACA by letting it perturb input nodes. Finally, the averaged length of the shortest paths over all environmental conditions and source attractors is reported. Formally, the average control path length determined by CABEAN for a BN model $M$, denoted $ACPL_{CABEAN}(M)$, id defined as follows:
$\text{ACPL}_{\text{CABEAN}}(M)\stackrel{{\scriptstyle\text{def}}}{{=}}\mathbb{E}_{e\in E_{M}}\left[\mathbb{E}_{s\in\mathcal{S}^{M}_{e}}\left[\begin{cases}\min\limits_{t\in\mathcal{T}^{M}_{e}}\ \text{cpl}(s,t)&\text{if }\mathcal{T}^{M}_{e}\neq\emptyset\\ \min\limits_{e^{\prime}\in E_{M}}\left(1+\min\limits_{t\in\mathcal{T}^{M}_{e^{\prime}}}\ \text{cpl}(s,t)\right)&\text{otherwise}\end{cases}\right]\right],$ (3)
where $E_{M}$ is the set of all environmental conditions considered for model $M$, $\mathcal{S}^{M}_{e}$ is the set of source attractors and $\mathcal{T}^{M}_{e}$ is the set of target attractors for a given environmental condition $e$ for model $M$, and cpl is the control path length computed by CABEAN for a specific pair of source and target attractors. Notice that in the case of a partially specified environmental condition where all the corresponding fully specified conditions can be handled by CABEAN at once, the set $E_{M}$ in Equation 3 is a singleton. In the case where all environmental conditions are considered, the cardinality of $E_{M}$ is equal to the number of distinct possible settings of the values for all input nodes.
6 Results
We evaluate GATTACA on the BN model environments presented in Section 5.1. The results obtained demonstrate that our approach is successful in controlling environments across a wide range of sizes.
6.1 Evaluation of iPASIP
Our iPASIP method, although it is a purely statistical method based on model simulations, provides a strong alternative to traditional methods for finding BN attractor states. Table 3 presents the summary results for each model where CABEAN was able to compute the attractors for the environmental conditions specified in Table 2.
To provide a more comprehensive evaluation of iPASIP, we additionally ran it on the Bladder, MAPK, T-LGL, and Bortezomib models without specifying the environmental condition – that is, simultaneously considering all of them. For comparison, we computed the attractors for each model with CABEAN in all individual environmental conditions one-by-one, i.e. in 16, 16, 128, and 32 conditions, respectively. The numbers of attractors in all environmental conditions are presented in square brackets in Table 3.
Model Size (Inputs) iPASIP PA States CABEAN
F-A C-A Spurious F-A C-A Failed
Bladder 35 (4) 9 [20] 0 [5] 0 [0] 9 [20] 0 [4] 1 of 16
MAPK 53 (4) 6 [20] 0 [0] 0 [0] 6 [20] 0 [0] 0 of 16
T-LGL 61 (7) 149 182 2? 166 130 6 of 128
Bortezomib 67 (5) 5 [83] 0 [0] 0 [0] 5 [83] 0 [0] 0 of 32
T-diff 68 (24) 12 0 0 12 0 0 of 1
MCF-7 117 (8) 4 0 0 4 0 N/A
CD4+ 188 (34) 12 0 1 12 0 0 of 1
Table 3: Comparison of the numbers of pseudo-attractor (PA) states identified by iPASIP with the numbers of attractors obtained with CABEAN. The numbers of fixpoint attractors (F-A), cyclic or complex attractors (C-A), and spurious pseudo-attractor states are presented for the environmental conditions in Table 2 (plain numbers) and for input nodes left unspecified, i.e. all possible environmental conditions (numbers in square brackets). The numbers of environmental conditions for which CABEAN failed to compute a result are shown in the ‘Failed’ column. The ‘?’ indicates that our verification of whether the pseudo-attractor states are indeed spurious failed.
For the Bladder model, iPASIP found 25 pseudo-attractor states in total in all the 16 environmental conditions, while CABEAN returned 24 attractors in 15 environmental conditions and failed in one, namely where the EGFR_stimulus input node is set to 1 and all the remaining three input nodes are set to 0. All the 24 CABEAN-found attractors are represented within the 25 pseudo-attractor states. By computing an attractor reachable from the extra pseudo-attractor state identified by iPASIP, we found that it is in fact an attractor state of an attractor consisting of 184,320 states under the environmental condition that CABEAN was unable to handle. We further investigated the distribution of revisits to the individual attractor states. We simulated a trajectory of 1,000,000 steps with the pseudo-attractor state as the initial state. The numbers of revisits for individual states are shown in Figure 3. The extra pseudo-attractor state is clearly dominating all other attractor states in terms of the count of revisits and is correctly identified by the iPASIP method despite the attractor being unusually large.
Figure 3: Bar plot of the number of revisits to individual states of the Bladder BN model attractor containing the extra pseudo-attractor state identified by iPASIP under the environmental condition of EGFR_stimulus set to 1 and all remaining input nodes set to 0. For readability, only states with at least 1,000 revisits are shown in the decreasing order of revisit count. State of index 0 is the pseudo-attractor state found by iPASIP and the remaining attractor states are indexed arbitrarily.
In two cases of T-LGL and CD4+, iPASIP overestimates the number of attractor states, i.e. some of the pseudo-attractor states are spurious attractor states, although for the T-LGL model we were unable to confirm that indeed these states are spurious: we failed to compute attractors reachable from the potentially spurious states with a graph traversal approach. Nevertheless, as argued in Section 4.1, the additional states stand out due to being frequently revisited. Thus, they can still be considered viable for performing control actions. Indeed, the extra pseudo-attractor state [^1] In the spurious attractor state the nodes Cofilin, GATA3, IL4, PI3K, proliferation, AKT, PIP3_345, Dec2, IRF4, IL4RA, PDK1, and IKB are 1 and all the remaining nodes are 0. of the CD4+ BN model is spurious – this can be inferred from the fact that CABEAN determined all attractor states for the given environmental condition. By simulating 1,000,000-step trajectories from this spurious attractor state, we observed that it was revisited more than 996,000 times and the trajectories consisted of less than 70 distinct states. Clearly, in this sense, the state resembles a stable state.
We verified that for all models except T-LGL all the attractors found by CABEAN have a representative amongst the pseudo-attractor states. For T-LGL, 143 of 166 CABEAN fixpoint attractors and 71 out of 130 CABEAN cyclic or complex attractors are also detected by iPASIP, some of them multiple times as more than one pseudo-attractor state may belong to the same attractor (data not shown). The remaining fixpoint pseudo-attractor states found by iPASIP are attractor states in the six environmental conditions in which CABEAN failed [^2] By failure we mean the fact that CABEAN ended with an error..
In summary, the iPASIP approach provides a scalable and effective solution for identifying true pseudo-attractor states, as defined in Definition 2.1, in large-size BN models. Moreover, as the case studies show, iPASIP often successfully identifies all attractor states of a large BN model.
6.2 Evaluation of the GATTACA control framework
With regard to control problem solutions found by GATTACA, the average length of the control strategies is typically within one control action/intervention of the optimal solutions computed in accordance with the formula in Equation 3 based on the control path lengths obtained by the exact ASI algorithm implemented in the CABEAN software tool, as shown in Table 4.
However, for the ABA model the number of environmental conditions is too big ($2^{23}$) to be handled by CABEAN (entries marked as N/A in Table 4). As computed by GATTACA, the ABA model is characterised by a large number of pseudo-attractor states. Nevertheless, GATTACA successfully identifies effective control strategies, demonstrating the scalability of our framework in this respect.
Model Size (Inputs) PA states Attractors GATTACA CABEAN
Bladder 35 (4) 9 9 1.0 1.0
MAPK 53 (4) 6 6 1.0 1.0
T-LGL 61 (7) 333 296 1.02 1.0
Bortezomib 67 (5) 5 5 1.99 1.05
T-diff 68 (24) 12 12 2.55 1.18
ABA 81 (23) 40,000 N/A 2.23 N/A
MCF-7 117 (8) 4 4 2.33 ?
CD4+ 188 (34) 13 12 1.16 1.03
Random200 200 (0) 165 N/A 7.1 N/A
Table 4: Comparison of the average control strategy lengths required to drive the network from the source state to the target configuration using GATTACA and the exact ASI method implemented in CABEAN.
To further investigate the scalability potential of GATTACA, we analyse the MCF-7 model under all possible environmental conditions. As in the case of the ABA model, MCF-7 is characterised by a large number of pseudo-attractor states and GATTACA is capable of successfully identifying control strategies. Unfortunately, we were unable to obtain the exact results with CABEAN as it could not finish the computations of attractors within one week time on an HPC cluster node where all $256$ fully specified environmental conditions were considered.
Next, we analysed the original T-diff model with 71 nodes, introduced in [20], under two environmental conditions: 1) the original, partially specified (PS) condition from [20] and 2) the fully specified (FS) condition from [29] presented in Table 2. The former condition leaves some of the input nodes unspecified with respect to the latter, i.e. IFNGR1, IFNGR2, GP130, IL6RA, CGC, IL12RB2, IL10RB, IL10RA, IL15RA, IL2RB, and IL27RA. We refer to these conditions as the partially and fully specified environmental conditions, respectively.
The original 71-node T-diff model, under both conditions, is characterised by a large number of pseudo-attractor states identified by iPASIP, i.e. 156,864 and 25,856 states, respectively. CABEAN fails to identify the attractors and cannot compute control strategies for these BNs. However, GATTACA successfully identifies control strategies in both cases. The average control strategy length is 6.17, as shown in Table 5. It is noticeably larger than for all other BN models considered in this study. Nevertheless, this is justifiable, as such large numbers of pseudo-attractor states suggest that the two BN models have large numbers of attractors. Meanwhile, the size of the original T-diff model (71 nodes) is comparable to the size of the model in [29] (68 nodes), the latter with only 12 attractors; see Table 4. It is then highly probable that the state space of the original T-diff model under the two environmental conditions is divided into much finer-grained basins of attraction than in the case of the modified model in [29]. Consequently, driving the dynamics towards a pseudo-attractor state aligned with the given target configuration requires an action that places the network dynamics into a relatively small basin of attraction, likely necessitating simultaneous perturbation of many genes. Since this is restricted due to the limit imposed on the maximum number of genes that can be perturbed at once within a single intervention (ATGC action), identified control strategies must traverse multiple attractors to reach the target configuration.
Model Size PA states GATTACA
MCF-7 117 1,500 2.11
Original T-diff in PS env. cond. 71 156,864 6.17
Original T-diff in FS env. cond. 71 25,856 6.17
Table 5: Average control strategy lengths obtained with GATTACA (shown in the ‘GATTACA’ column) for the original 71-node T-diff model of [20] under two environmental conditions (env. cond.), i.e. the partially specified (PS) one in [20] and the fully specified (FS) one in [29]. Both conditions are characterised by large number of pseudo-attractor states identified by iPASIP.
Finally, we perform one additional scalability test of GATTACA on a randomly generated BN model of 200 nodes (Random200). The model is challenging as it has no input nodes and its structure graph consists of a single SCC. Therefore, the attractor detection algorithm and the ASI control algorithm in CABEAN cannot exploit the SCC-decomposition and CABEAN fails to compute the control strategies within 168 hours (one week) on a machine with 8 Core 2.25 GHz AMD EPYC 7742 processor and 256 GB of RAM. GATTACA identifies 165 pseudo-attractor states and returns control strategies of average length of 7.1 for the target configuration $x_{3}=1$, where $x_{3}$ is one of the nodes; see Table 4. This relatively long control strategy on average can be explained by the fact that a network of highly interconnected nodes may require a simultaneous perturbation of a larger number of nodes than in the case of a loosely connected network. Similarly to the two cases of the original T-diff model in Table 5, since we limit the number of simultaneous perturbations, GATTACA finds longer control strategies that traverse a number of intermediate pseudo-attractor states.
7 Conclusions
We formulated the attractor-target control problem in the context of cellular reprogramming –a novel control problem for biological networks modelled using the BN framework under the asynchronous update mode. To provide a scalable solution, we developed and implement GATTACA, a scalable computational framework based on DRL.
To facilitate scalability of GATTACA, we consider our previously introduced concept of a pseudo-attractor and we further develop our PASIP procedure for effective identification of pseudo-attractor states by properly handling the situations in which model simulations get trapped in large attractors. We demonstrate the effectiveness of improved PASIP, i.e. iPASIP, on a number of BN models of real-world biological networks of different sizes by comparing the identified pseudo-attractor states with true attractors obtained with an exact state-of-the-art algorithm implemented in the CABEAN software tool. We show that iPASIP is capable of correctly identifying representative attractor states of all attractors in most of the cases. It also manages to detect attractor states in BN models for which CABEAN fails to compute the attractors. Furthermore, iPASIP proves its scalability and effectiveness as part of the GATTACA framework, where the determined pseudo-attractor states facilitate the computation of control strategies for large BN models of various real-world biological networks.
The proposed GATTACA framework offers a scalable and effective solution to the attractor-target control problem. To leverage the known structure of biological systems for encoding the network dynamics in a latent representation, we incorporate graph neural networks with graph convolutions into the artificial neural network approximator for the action-value function learned by the ATGC DRL agent. Given source attractor state and target configuration of genes, the ATGC of GATTACA finds proper control strategies that drive the network from the source pseudo-attractor to the target configuration by taking actions (intervening) only in intermediate pseudo-attractor states that correspond to phenotypical cellular states that can be observed in the lab.
The results of the conducted computational experiments on a number of large real-world biological networks from literature demonstrate the scalability and effectiveness of GATTACA. Our framework is capable of finding optimal or suboptimal control strategies which lengths in most of the cases differ on average from the optimal strategies by at most one intervention. Moreover, GATTACA is capable of identifying effective control strategies of reasonable lengths in BN models on which CABEAN fails. These models are large in terms of number of nodes and often characterised by many (pseudo-)attractor states.
An additional evaluation on a challenging network which structure graph is a single large SCC of 200 nodes further demonstrates the scalability potential of GATTACA in controlling large models of complex biological networks.
8 Acknowledgments
This research was funded in whole or in part by the National Science Centre, Poland under the OPUS call in the Weave programme, grant number 2023/51/I/ST6/02864. For the purpose of Open Access, the author has applied a CC-BY public copyright licence to any Author Accepted Manuscript (AAM) version arising from this submission.
Appendix A Example
(0,0,0)(0,0,1)(0,1,0)(0,1,1)(1,0,0)(1,0,1)(1,1,0)(1,1,1)
Figure 4: Example of an STG
Let us consider a Boolean Network defined as:
$\begin{cases}x_{1}=x_{1}\\ x_{2}=x_{1}|x_{3}\\ x_{3}=x_{2}\&x_{3}\end{cases}$
It’s STG is shown in the Pic. 4. In this network $x_{1}$ is an input node, that is its state remains constant during the evolution of the network. Let us consider a target of the form $x_{2}=0$. That is we want our network to be in an attractor with $x_{2}$ set to 0, and we do not care for the state of other nodes.
If we consider the case of $x_{1}=0$ then our target is only realized by $(0,0,0)$ state, and can be controlled from the other attractor by one perturbation:
[…]
In the case of $x_{1}=1$ we have one state ralizing the target: $(1,0,0)$. It will naturally be reached via evolution from $(1,1,0)$ and we will need one perturbation to reach it from $(1,1,1)$:
[…]
Appendix B Evaluation example
We consider a BN with 5 nodes, i.e. $x_{1},x_{2},x_{3},x_{4},x_{5}$. The nodes $x_{1}$ and $x_{2}$ are designated as input nodes and the control objective is to achieve the target configuration $x_{5}=1$. Let us assume that
  • For $x_{1}=0,x_{2}=*$, the network has two attractors:
    $(0,0,0,0,0),\quad(0,0,1,1,0)$
    In both cases $x_{5}=0$, so there is no attractor aligned with the target configuration.
  • For $x_{1}=1,x_{2}=1$, the attractors are:
    $(1,1,1,0,1),\quad(1,1,0,1,0)$
  • For $x_{1}=1,x_{2}=0$, the attractor is:
    $(1,0,0,0,1)$
Thus, there are three source attractors and two target attractors.
We consider $(0,0,0,0,0)$ to be the source attractor. Without perturbations of the input nodes $x_{1}$ and $x_{2}$, the system remains trapped in the configuration $x_{5}=0$. Therefore, perturbations of input nodes are necessary to change the environmental condition. There are two options: set $x_{1}=1,x_{2}=1$ or $x_{1}=1,x_{2}=0$. We consider both.
  • Suppose that the intervention to state $(1,1,0,0,0)$ leads to attractor:
    $(1,1,0,1,0)$
    Since $x_{5}=0$, an additional intervention on $x_{3}$ is required to reach the target attractor, which is a CABEAN control path of length 1. Therefore, the total control strategy length in this case is 2.
  • The intervention to state $(1,0,0,0,0)$ leads directly to attractor:
    $(1,0,0,0,1),$
    which is a target attractor. Here, the total control strategy length is 1.
Then, the length of the reference control strategy obtained with CABEAN is 1 for this example.